OpenStudy (anonymous):

how do you go from (2cos^2x+2sinx+2sin^2x)/cos^2 to (2+2sinx)/cos^2x

OpenStudy (anonymous):

2cos^2x + 2sin^2x = 2(cos^2x + sin^2x) = 2(1) =2